OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 1 < 9 or 2x >24 - - A number line with an open circle on 10, shading to the left, and an open circle on 12, shading to the right. A number line with an open circle on 10, an open circle on 12, and shading in between. A number line with a closed circle on 10, a closed circle on 12, and shading in between. A number line with a closed circle on 10, shading to the left, and a closed circle on 12, shading to the right.

OpenStudy (anonymous):

\[x - 1 \le 9 \]+1\[x \le 10\] \[2x \ge 24\]/2\[x \ge 12\]

OpenStudy (anonymous):

So they will both be closed circles. The key is that it says "or" so the answer would be D.A number line with a closed circle on 10, shading to the left, and a closed circle on 12, shading to the right.

OpenStudy (anonymous):

please continue to help me :) thanks :)

OpenStudy (anonymous):

k, np:)